OpenStudy (anonymous):

Someone Please help me! Suppose you have a 30-year home loan with an interest rate of r percent, compounded monthly.The amount you borrow is P dollars. How much is the monthly payment M?

OpenStudy (dumbcow):

did you just want the formula? \[P = M (\frac{v(1-v^{360})}{1-v})\] where \[v = \frac{1}{1+r}\] alternative formula can be found here http://www.financeformulas.net/Loan_Payment_Formula.html